OpenStudy (anonymous):

Denny, Jamal, and Bobby are saving up quarters. Denny has one fifth as many as Bobby, and Jamal has three fewer than twice as many as Denny. All three of them combined have 61 quarters. Find how many each has. (there was another one of these on Google search, but it didn't really have it solved as a single equation)

OpenStudy (texaschic101):

D + J + B = 61 D = 1/5B J = 2(1/5B) - 3 now its just a matter of subbing.. (1/5B) + 2(1/5B) - 3 + B = 61 1/5B + 2/5B - 3 + B = 61 1/5B + 2/5B + 5/5B = 61 + 3 8/5B = 64 B = 64/(8/5) B = 64 * 5/8 B = 40 D = 1/5B D = 1/5(40) D = 8 J = 2(1/5(40) - 3 J = 2(8) - 3 J = 16 - 3 J = 13 check.. D + J + B = 61 8 + 13 + 40 = 61 8 + 53 = 61 61 = 61 (correct)

OpenStudy (anonymous):

thank you! the fraction was throwing me off and I also forgot to add Bobby's amount by itself to the equation.